find the volume of a cone when the radius =1 1/2 and the height = 21 yards
Answer by binayakojha(1)

here,
radius (r)= 1 1/2=3/2=1.5
height (h)= 21
volume of cone = (1/3)* 3.14*r^2*h
=(1/3)* 3.14*(1.5)^2*21
= 49.455 cubic yard